You are the buyer for your university bookstore. One of the textbooks has a cost to you of $155 and you sell it to students for $310. In this case, however, you cannot salvage any value from copies that do not sell because a new edition is published every semester. Demand for this text averages 135 copies each semester, with a standard deviation of 21 copies. Use Appendix A. How many should you order each semester? (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole number.)
Answer #1

Purchase Price = $155

Selling Price = $310

Salvage Price = $0

Cost of Underselling Cu = Purchase price - Salvage Price = 155 - 0 = 155

Cost of Overselling Co = Selling Price - Purchase Price = 310 - 155 = 155

p <= Cu/(Cu+Co) = 155/310 = 0.50

From z table, for p = 0.50, z = 0

Let the number of books to be ordered be X

Average Demand = μ = 135

Standard Deviation = σ = 21

Hence, z = (X - μ )/ σ

=> 0 = (X - 135 ) / 21

=> X = 135

Hence, you should order 135 copies each semester